# Extended Resolution Clause Learning via Dual Implication Points

### v1

@@ 6.2 Performance Analysis on 2023 SAT Competition Benchmarks
Our experimental evaluation concludes with the lessons we have learned from executing our DIP-variants [^4] We do not present a comparison with state-of-the-art solvers like Kissat or CaDiCaL since there are not many lessons to be learned. In a few words, our MapleLCM baseline CDCL solver is slower that those systems, and adding DIP-reasoning to it does not bridge this gap. on all benchmarks from the 2023 SAT Competition. We first report on the overhead caused by the DIP detection algorithm and the subsequent additional work to retrieve the clauses to be learned. For our baseline DIP-based system, where two clauses are learned and hence is the most computationally demanding method, in 4.5% of the benchmarks the DIP-related work represented between 10 and 15% of the total runtime; in 11.5% of the benchmarks between 5 and 10%; in 35.5% between 2.5 and in 48.5% less than 2.5%. These data show that DIP computation does not significantly slow down the solver. Another interesting information concerns the percentage of conflicts where there is at least one DIP, which was on average 63%. This implies that we do need to have filtering mechanisms to discard some of them. Otherwise, the search would be totally dominated by DIPs.
A cactus plot with the results of executing all our DIP variants on the 2023 SAT Competition benchmarks with a time limit of one hour is in the left plot of Figure 8. Three variants are much worse than the rest: act, occ5 and glue, whereas the other are relatively similar. These three configurations have a much larger percentage of conflicts where DIP-based learning is applied than the rest, being around 30% in act and around 5% in the other two. In the best performing variant this is around 1%, which seems to be a very low number and one might wonder whether this is enough to have any effect at all. This can be answered by analyzing the percentage of decisions on extended variables. For all benchmarks where DIP-based systems outperform the baseline CDCL solver, this number is surprisingly large, being usually over 10% of the decisions, whereas for most of the benchmarks is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables). Moreover, this does not only happen in problems with a very low number of initial variables, where the introduction of a few extended variables could quickly dominate the decision heuristic.
@@
This is remarkable since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system. This is why we implemented, on top of our baseline DIP system, a procedure that computes the percentage of decisions on extended variables. If after a given number of conflicts it is still lower than 3%, it discontinues DIP learning and performs 1-UIP learning from that moment onwards. This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found in the right plot of Figure 8, is very positive.
Figure 8: Performance of DIP variants on 2023 SAT Competition benchmarks.
7 Conclusions and Future Work
@@
We introduce a novel extended resolution clause learning (ERCL) algorithm, that when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random k-xor and interval matching formulas. This is remarkable since automating powerful proof systems is well-known to be a difficult task. Somewhat surprisingly, the only previously existing attempt to incorporate ER into CDCL performs similarly on the same instances. Considering the different nature of the two methods, this clearly deserves further study. Further, we also introduce a new heuristic that allows our ERCL solver xMapleSAT to perform similarly to the baseline CDCL solver MapleLCM, thus being able to get the best of both worlds, i.e., the benefit of ERCL, without sacrificing performance of the CDCL solver on, say, SAT Competition Main Track 2023 instances. As future work, we plan to investigate a variety of machine learning based heuristics (e.g., branching) specialized for the ERCL method. Also, the use of $k$-IPs with $k>2$ is part of our next steps. On the theoretical side, challenging questions like determining whether DIP- or $k$-IP based ER simulates unrestricted ER are going to be central to our research efforts.

### v2

@@ 6.4. Performance Analysis on SAT Competition Benchmarks
Our experimental evaluation concludes with the lessons we have learned from executing our DIP implementation on the SAT competition benchmarks. We used benchmarks from the 2023 and 2024 editions. In the latter, the number of instances from the Tseitin, Xorified random $k$-xor and Interval families we have described is much lower than in the 2023 edition, and hence we present results from the 2023 competition in order to show the possible gains or losses of adding DIP-based reasoning to solvers. We first report on the overhead caused by the DIP detection algorithm and the subsequent additional work to retrieve the clauses to be learned. For our baseline DIP-based system, where two clauses are learned and hence is the most computationally demanding method, in 4.5% of the benchmarks the DIP-related work represented between 10 and 15% of the total runtime; in 11.5% of the benchmarks between 5 and 10%; in 35.5% between 2.5 and in 48.5% less than 2.5%. These data show that DIP computation does not significantly slow down the solver. Another interesting information concerns the percentage of conflicts where there is at least one DIP, which was on average 63%. This implies that we do need to have filtering mechanisms to discard some of them. Otherwise, the search would be totally dominated by DIPs.
In the left cactus plot of Figure 9 we can see that our baseline DIP-based xMapleLCM solver outperforms the CDCL-based MapleLCM that it is based on. When we analyzed the concrete benchmarks where our DIP-based solver outperforms MapleLCM and other CDCL-based solvers like Kissat, we observed that for all of them the percentage of decisions on extended variables is surprisingly large (over 10% of the decisions), whereas for the remaining benchmarks it is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables). Moreover, this happens no matter how many initial variables there are, even though in problems with few number of initial variables the introduction of a few extended variables could quickly dominate the decision heuristic.
@@
This is remarkable since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system. This is why we implemented, on top of our baseline DIP system, a procedure that computes the percentage of decisions on extended variables. If after a given number of conflicts it is still lower than 3%, it discontinues DIP learning and performs 1-UIP learning from that moment onwards. This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found on the left cactus plot of Figure 9, is very positive, showing that a solver disabling DIP reasoning (xMapleLCM-disabling in the plot) when the percentage of decisions on extended variables is low is still able to solve Tseitin, random $k$-xor and Interval benchmarks, but the overall performance is not significantly worsened on the remaining instances. Finally, the right cactus plot shows the impact of changing the type of DIP into the disabling solver: the heuristic and middle DIPs, the latter being used in the baseline configuration, are much better than random and closest ones.
Figure 9. Performance of DIP variants on 2023 SAT Competition benchmarks.
Learned lessons from experimental evaluation. Let us finish this section by pointing out the main lessons we have learned from this experimental evaluation.
  • We have identified three families of benchmarks for which the dynamic addition of extension variables improves upon CDCL in a systematic manner. This is an improvement w.r.t. the results in the GlucosER paper [AKS10], where a few initial benchmarks were identified but it was not proved that extended-resolution methods always outperform CDCL on increasingly more complex instances.
  • We have proved that, even though the way extended variables are identified is completely different in xMapleLCM and GlucosER, they both behave well on the same instances. This is remarkable because our preliminary experiments revealed that randomly choosing arbitrary pairs of (non-DIP) literals to create definitions does not provide any speedup w.r.t. CDCL. Hence, both methods are able to identify semantic properties about concrete literals that make them suitable for creating an extended variable definition.
  • We have learned that in most of the families where extended resolution works there is a large xor component. Additionally, when this is xor information is somehow hidden in the formula, our dynamic approach outperforms static methods like the ones in CryptoMiniSat and SBVA-CaDiCaL.
  • Our ablation experiments have shown that the most important design choice is the one of determining which DIP is to be chosen. Among all the possibilities we have devised, middle and heuristic are the preferable ones.
  • Finally, we have noticed that by inspecting the percentage of decisions on extended variables, one can detect whether our DIP-based is helping the solver. This has allowed us to further improve the behavior of our solver on sets of benchmarks where only a few of them are advantageous for extended resolution.
7. Conclusions and Future Work
@@
We introduce a novel extended resolution clause learning (ERCL) algorithm that, when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random $k$-xor and interval matching formulas. This is remarkable since automating powerful proof systems is well-known to be a difficult task. Somewhat surprisingly, the only previously existing attempt to incorporate ER into CDCL performs reasonably well on the same instances. We have hence indentified classes of formulas for which extended resolution can be automated in at least two different ways and still outperform CDCL. Considering the different nature of the two methods, this clearly deserves further study. Further, we also introduce a new heuristic that allows our ERCL solver xMapleSAT to perform similarly to the baseline CDCL solver MapleLCM, thus being able to get the best of both worlds, i.e., the benefit of ERCL, without sacrificing performance of the CDCL solver on, say, SAT Competition Main Track 2023 instances.
As future work, we plan to investigate a variety of machine learning based heuristics (e.g., branching) specialized for the ERCL method. Also, the use of $k$-IPs with $k>2$ is part of our next steps. On the theoretical side, challenging questions like determining whether DIP- or $k$-IP based ER simulates unrestricted ER are going to be central to our research efforts.